# MultiGA: Leveraging Multi-Source Seeding in Genetic Algorithms

### v1

@@ Abstract
Large Language Models (LLMs) are widely used across various research domains to tackle complex tasks, but their performance can vary significantly depending on the task at hand. Compared to fine-tuning, inference-time optimization methods offer a more cost-effective way to improve LLM output. Evolutionary algorithms can be used to refine solutions iteratively, mimicking natural selection. To the best of our knowledge, there has not been exploration on leveraging the collective capabilities of multi-source seeding for LLM-guided genetic algorithms. In this paper, we introduce a novel approach, MultiGA, which applies genetic algorithm principles to address complex natural language tasks and reasoning problems by sampling from a diverse population of LLMs to initialize the population. MultiGA generates a range of outputs from various parent LLMs, open source and closed source, and uses a neutral fitness function to evaluate them. Through an iterative recombination process, we mix and refine these generations until an optimal solution is achieved. Our results show that MultiGA converges to the accuracy of the LLM best fit for the task, and these insights lay the foundation for future research looking closer at integrating multiple LLMs for unexplored tasks in which selecting only one pre-trained model is unclear or suboptimal. We benchmark our approach using text-to-SQL code generation tasks, trip planning, GPQA benchmark for grad-level science questions, and the BBQ benchmark that measures bias in models. This work contributes to the growing intersection of evolutionary computation and natural language, highlighting the potential of biologically inspired algorithms to improve generative artificial intelligence selectivity and accuracy.
@@ 1 Introduction
Figure 1: Overview of the MultiGA framework. Populations are initialized with multiple LLMs, while an independent LLM $E$ handles fitness evaluation (scoring candidates) and recombination (combining two parent solutions). The process terminates once target fitness $\phi$ or maximum number $T$ generations is reached. Then, the top candidate solution is returned.
The development of small language models (SLMs) and pretrained language models (PLMs) marked early progress in natural language processing, opening new possibilities for text understanding and generation. Models such as BERT (devlin_bert_2019) and RoBERTa (liu_roberta_2019) demonstrated the power of large-scale pretraining, while ULMFiT (howard_universal_2018) showcased the utility of fine-tuning for downstream tasks. However, these models struggle with unfamiliar prompts and PLMs often require extensive task-specific engineering, which limits their applicability (zhu_impossible_2022). The emergence of large language models in 2020 marked a turning point: models such as GPT-3 demonstrated, for the first time, strong generalization across a wide range of tasks (brown_language_2020). These early LLMs achieved substantially higher accuracy than smaller pretrained models on closed-book question answering benchmarks like TriviaQA (joshi_triviaqa_2017), and showed significant gains in arithmetic reasoning and other challenging domains (brown_language_2020).
Furthermore, LLMs support one-shot and few-shot prompting, reducing the need for fine-tuning and making them attractive for workflows that require flexible reasoning (zhao_-depth_2023). Prompting techniques such as Chain-of-Thought (CoT) (wei_chain--thought_2023) and Tree-of-Thoughts (ToT) (yao_tree_2023) further enhance performance by structuring reasoning into sequential or branching steps, effectively breaking complex tasks into manageable components for the LLM to interpret. This idea of decomposing tasks has driven the rise of multi-agent workflows, particularly in industry applications. Consider text-to-SQL: solving a single query may require preprocessing natural language, linking question terms to database schema, generating SQL code, and validating outputs. Complex pipelines like this often adopt multiple agents, where each specialized LLM agent handles a subtask. Nevertheless, accuracy remains a persistent challenge, especially for developers who rely on open-source models or have limited access to task-optimized models. This challenge is especially evident when it is unclear which model is best suited for novel problems requiring interdisciplinary skills, a difficulty further compounded by the heterogeneity of LLMs, whose outputs vary significantly depending on their source.
@@
The quality and performance of LLMs vary based on how they are trained, leading to complementary strengths and weaknesses. Beyond large-scale pretraining, most modern systems incorporate instruction tuning and reinforcement learning from human feedback (RLHF) to align with user intent (ouyang_training_2022). Open families such as LLaMA apply instruction tuning and lightweight RLHF to achieve strong performance with smaller parameter budgets (touvron_llama_2023). Furthermore, more recent efforts, like DeepSeek R1, extend this paradigm with reinforcement learning from AI feedback (RLAIF) and prompting-based curricula to enable improved reasoning without extensive human annotation (deepseek-ai_deepseek-r1_2025). These design choices produce models with distinct advantages, pointing toward strategies that can exploit their complementarities.
@@
With the advent of this new generation of reasoning models, inference-time self-improvement has emerged as a popular application. Techniques such as Self-Refine allow LLMs to generate an initial answer, critique it, and iteratively refine the response, yielding better results without additional supervised training or RLHF (madaan_self-refine_2023). Self-reflection has also been used in agentic frameworks where agents enhance problem solving by analyzing incorrect outputs, diagnosing errors, and adjusting future reasoning paths (renze_self-reflection_2024). Furthermore, reflection has also been extended to the concept of LLMs-as-Judges, where one model critiques or evaluates the outputs of another (li_llms-as-judges_2024). This approach reduces reliance on costly human annotations and enables fine-grained feedback, making it a powerful tool in agentic settings.
@@
Taken together, the diversity in model training and inference strategies suggests that different LLMs contribute unique strengths. Incorporating multiple LLMs within the same framework creates a broader and more diverse solution space. Concepts from evolutionary computation provide a natural lens for exploring this diversity, as variation and recombination can be used to amplify strong solutions and suppress weaker ones. Building on this perspective, our key contributions in this paper are:
@@
1) MultiGA framework: We propose MultiGA, a genetic optimization framework that seeds the initial population with outputs from multiple LLMs and employs an independent evaluator LLM.
2) Empirical validation: We demonstrate that MultiGA reduces reliance on any single model by expanding the candidate pool and exploiting complementary models, offering a practical approach for achieving high accuracy across diverse tasks.
3) Foundation for future work: We establish a basis for exploring how ensembles of diverse strong models can further improve performance on interdisciplinary and novel domains.
@@ 2 Related Works
2.1 Evolutionary Algorithms
Evolutionary algorithms (EAs) draw inspiration from natural selection and have demonstrated strong applicability not only in machine learning but also across a wide range of academic disciplines (goldberg_genetic_1989). Candidate solutions are generated and refined through a process of evaluation, selection, and recombination, including crossover and mutation. Typically, the initial population is generated at random, and each solution is evaluated using a fitness function to determine which candidates are retained and combined with others to produce new “children”. The appeal of using these algorithms lies in the ability to efficiently navigate large, complex search spaces and identify high-quality solutions, rather than being limited to a single output. As newer generations of solutions are added into the population, older and less fit solutions are retired. Traditionally, evolutionary algorithms employ predefined rules for crossover and mutation—often arbitrary or randomized—and rely on fixed algorithmic fitness functions to evaluate candidate solutions.
@@ 2.2 Genetic Algorithms and Programs
Genetic algorithms (GA), which usually operate on simpler fixed-length encodings like bitstrings or arrays, are often used for optimization tasks where the solution can be encoded as a sequence of values. GAs search the solution space efficiently and have been widely applied in domains such as reinforcement learning (whiteson_evolutionary_2006) and prompt optimization (secheresse_gaapo_2025). Genetic programming (GP), which evolves tree-structured representations, has specifically been used for Automatic Heuristic Design to automatically generate and adapt decision-making rules tailored to specific tasks (burke_automatic_2007). These processes have advanced with the introduction of LLM-based evolutionary program search, where large language models guide the exploration, refinement, and assessment of complex programs in place of these less generalizable, rule-based frameworks (zhang_understanding_2024; shum_hybrid_2025).
2.3 LLM-Guided Evolutionary Algorithms
Recent work has explored the intersection of evolutionary algorithms and large language models in several ways. LLMs have been employed as mutation operators within evolutionary frameworks, leveraging their semantic understanding to generate more meaningful variations (lehman_evolution_2022). Evolutionary approaches have also been applied to neural architecture search, automatically discovering optimal network structures (wistuba_survey_2019). In the domain of automated algorithm design, Quality-Diversity algorithms such as MAP-Elites have been used to discover diverse collections of high-performing solutions (mouret_illuminating_2015). Furthermore, evolutionary prompt engineering has emerged as a method to optimize prompts for specific tasks, treating prompts as evolvable entities subject to selection and variation (zhou_large_2023). These approaches collectively demonstrate the potential of combining evolutionary computation with modern language models to automate and enhance the design of algorithms, heuristics, and prompts.
@@ 3 Methodology
3.1 Overview
We introduce the Multi-Source Genetic Algorithm a framework designed to harness strengths from a diverse pool of LLMs. Rather than relying on a single model for both generation and evaluation, MultiGA seeds its initial population with outputs from multiple LLMs, each bringing unique variations shaped by its training and architecture. These candidate solutions are then iteratively recombined and refined through an independent evaluator model, which provides unbiased feedback and guides the selection of stronger offspring. MultiGA aims to create a search process that converges toward high-quality solutions while preserving diversity in the solution space, which will be explained in the following sections.
@@ 3.4 Evaluation and Selection
In order to guide the search toward progressively more accurate solutions, we apply a fitness evaluation to every candidate in the population at each iteration of the algorithm. At generation $t$, the population is represented as $P_{t}=\{x_{1},x_{2},\dots,x_{n}\}$. Each candidate $x_{i}$ is assigned a fitness score by $E$, defined as
@@
where $s_{i}=1$ corresponds to a perfect solution and $s_{i}=0$ indicates an invalid one. The evaluator $E$ is provided with all relevant task information (e.g., the query, context, or constraints) and is instructed to judge the overall correctness and quality of the candidate. This approach ensures that scoring is based on semantic adequacy rather than superficial string similarity.
@@ 3.5 Recombination
To generate new candidates, we employ an LLM-based crossover operator using the same independent LLM $E$. Given two parent solutions, $E$ is provided with all task information along with the parent solutions and is prompted to synthesize a child that integrates the strengths of both. Unlike traditional token-level crossover rules, this operator leverages the generative capacity of LLMs to perform semantic recombination. The resulting offspring are diverse yet coherent, inheriting useful attributes from each parent while overcoming the constraints of rule-based recombination strategies.
Figure 2: Example of recombination on a text-to-SQL task (“Which publisher published the slowest superhero?”). Parent 1 (gpt-4o, score 0.95) was paired with a randomly selected parent (deepseek-r1, score 0.1). The resulting child achieved a perfect score of 1.0 by preserving gpt-4o’s overall structure while incorporating the MIN aggregation from r1.
@@ 4 Experiments
4.1 Overview
To assess whether seeding genetic algorithms with multiple LLMs enables deeper exploration, we evaluate our framework across four tasks. Our aim is not to design highly specialized or task-specific algorithms, but rather to provide a simple and general framework in which all models can be evaluated fairly. While previous GA-based methods like Mind Evolution (lee_evolving_2025) have demonstrated promising results, challenges in reproducibility motivate our focus on developing an accessible and well-documented implementation.
@@ 4.1 Overview
In all experiments, we used a consistent configuration across tasks to ensure comparability. We selected the top-3 ($k=3$) parents for recombination, each run was capped at a maximum of $T=3$ generations, we used a fitness threshold of $\tau=0.2$ for pruning low-quality candidates, and our target fitness was $\phi=0.95$ as the stopping criterion. These conservative settings were chosen to keep the computational cost manageable while still allowing the algorithm to demonstrate meaningful improvements over successive generations. The genetic algorithm itself is implemented as a general class that can be configured for any task using prompts defined in a task-specific configuration file. For all tasks, we designate gpt-4o-mini as the unbiased evaluator and recombination model, responsible for scoring candidate solutions and synthesizing offspring across generations. While the choice of evaluator was somewhat arbitrary, we required a model strong enough to serve as a reliable judge. Notably, our results indicate that the use of gpt-4o-mini did not unduly inject its own expertise, as certain models still performed poorly on the benchmarks despite its assessment and recombination guidance.
@@
$\mathcal{D}_{\texttt{SQL}}$ (BIRD Mini-Dev) 100 500 $\infty$
@@
$\mathcal{D}_{\texttt{BBQ}}$ (BBQ Bias Data) 104 6879 3
@@
Table 1: Datasets used in our experiments, with test set size, total available items, and number of label choices.
4.2 Baselines
As a point of comparison, we evaluate MultiGA against single-LLM baselines, where the genetic algorithm design remains identical but the initial population is seeded exclusively from one model. The individual LLMs we use are:gpt-4o (OpenAI), qwen2.5-coder-32b-instruct (Qwen), deepseek-r1-distill-qwen-32b (DeepSeek), openai/gpt-oss-20b (OpenAI OSS), and mistral-7b-instruct-v0.3 (Mistral). These models were chosen for their accessibility and and provide a diverse set of baselines for comparison.
The “ALL” condition represents the full MultiGA framework, in which the initial population is seeded with outputs from all five models simultaneously. This design isolates the impact of diverse initialization while keeping all other components of the algorithm fixed. For comparison, we also evaluated the same population initialization prompts without applying the iterative refinement of MultiGA.

### v2

@@ Abstract
In this paper, we introduce, MultiGA, an optimization framework which applies genetic algorithm principles to address complex natural language tasks and reasoning problems by sampling from a diverse population of LLMs to initialize the population of candidate solutions. MultiGA generates a range of outputs from various parent LLMs and uses a neutral fitness function to evaluate them. Through an iterative recombination process, we mix and refine these generations until an optimal solution is achieved. Our results show that MultiGA produces high accuracy across multiple benchmarks, and these insights lay the foundation for future research looking closer at integrating multiple LLMs for unexplored tasks in which selecting only one pre-trained model is unclear or suboptimal.
@@ 1 Introduction
Figure 1: Overview of the MultiGA framework. Populations are initialized with multiple LLMs, while an independent LLM $E$ handles fitness evaluation (scoring candidates) and recombination (combining two parent solutions). The process terminates once target fitness $\phi$ or maximum number $T$ generations is reached. Finally, through iteration, the top candidate solution is evolves with the feedback and knowledge mixing from other LLMs.
Over the past few years, research in Large Language Models (LLMs) have radically transformed human workflows, particularly in supporting problem-solving across diverse domains. These applications extend beyond computational tasks, demonstrating effectiveness in specialized disciplines such as healthcare, law, and the natural sciences (tran2025multiagent). While LLMs can exploit a large corpora of human knowledge, their reasoning capabilities remain uneven, with notable performance disparities across proprietary models (e.g., Gemini, GPT, Claude), as reflected in their performance across niche reasoning benchmarks (livebench). In our research, we use the MultiGA framework to explore the complementary strengths of various LLMs to enhance performance and robustness on reasoning-intensive tasks while reducing reliance on any single model family.
@@
The quality and performance of LLMs vary based on how they are trained, leading to complementary strengths and weaknesses. Beyond large-scale pretraining, most modern systems incorporate instruction tuning and reinforcement learning from human feedback (RLHF) to align with user intent (ouyang_training_2022). Open families such as LLaMA apply instruction tuning and lightweight RLHF to achieve strong performance with smaller parameter budgets (touvron_llama_2023). More recent efforts, like DeepSeek R1, extend this paradigm with reinforcement learning from AI feedback (RLAIF) and prompting-based curricula to enable improved reasoning without extensive human annotation (deepseek-ai_deepseek-r1_2025). Additionally, the Alibaba Qwen3 series was trained across 119 languages and utilizes a post-training process that integrates both "thinking" and "non-thinking" modes, allowing a single model to switch between complex multi-step reasoning and direct, context-based responses (yang_qwen3_2025). These design choices produce models with distinct advantages, pointing toward strategies that can exploit their complementarities.
Prompting techniques such as Chain-of-Thought (CoT) (wei_chain--thought_2023) and Tree-of-Thoughts (ToT) (yao_tree_2023) further enhance performance by structuring reasoning into sequential or branching steps. This idea of decomposing tasks has driven the rise of multi-agent workflows, particularly in industry applications. Consider text-to-SQL: solving a single query may require pre-processing natural language, linking question terms to database schema, generating SQL code, and validating outputs on the database. Pipelines like this often adopt multiple agents, where each specialized LLM agent handles a sub-task. Nevertheless, accuracy remains a challenge, especially for developers who rely on open-source models or have limited access to task-optimized models. This challenge is especially evident when it is unclear which model is best suited for new problems requiring interdisciplinary skills.
@@
With this new generation of reasoning models and prompting techniques, inference-time self-improvement has emerged as a popular tool. Techniques such as Self-Refine allow LLMs to generate an initial answer, critique it, and iteratively refine the response, yielding better results without additional supervised training (madaan_self-refine_2023). Self-reflection has also been used in agentic frameworks where agents enhance problem solving by analyzing incorrect outputs, diagnosing errors, and adjusting future reasoning paths (renze_self-reflection_2024). Furthermore, reflection has also been extended to the concept of LLMs-as-Judges, where one model critiques or evaluates the outputs of another (li_llms-as-judges_2024).
@@
Taken together, the diversity in model training and inference-time behavior suggests that different LLMs contribute unique strengths. Incorporating multiple LLMs within the same framework creates a broader and more varied solution space. Concepts from evolution, provide a natural lens for exploring this diversity, as processes based on natural selection can be used to amplify strong solutions and suppress weaker ones. Our key contributions in this paper are:
@@
1) MultiGA framework: We propose MultiGA, a genetic optimization framework that seeds the initial population with outputs from multiple LLMs and employs an evaluator and recombination LLM to refine the candidate solutions until an optimal solution is achieved.
2) Empirical validation: MultiGA reduces reliance on any single model, achieving strong accuracy across varied tasks. On three of four benchmarks, MultiGA seeded with a primarily open-source ensemble outperforms a GPT-5-only seeded variant (both using the same evaluator), and surpasses multiple single-model baselines.
3) Foundation for future work: This research lays the groundwork for exploring how ensembles of heterogeneous, open-source and proprietary models can drive further gains across interdisciplinary and novel domains.
@@ 2 Related Works
2.1 Evolutionary Algorithm Overview
Evolutionary algorithms are search algorithms inspired by natural selection and have demonstrated strong applicability across a wide range of academic disciplines (goldberg_genetic_1989). Candidate solutions are generated and refined through a process of evaluation, selection, and recombination, including crossover and mutation. Traditionally, the initial population is generated at random, and each solution is evaluated using a fitness function to determine which candidates are retained and combined with others to produce new “children”. The appeal of using these algorithms lies in the ability to efficiently navigate large, complex search spaces and identify high-quality solutions, rather than being limited to a single output. As newer generations of solutions are added into the population, older and less fit solutions are retired. Evolutionary algorithms often employ predefined rules for crossover and mutation and rely on fixed algorithmic fitness functions to evaluate candidate solutions. Taken together, this framework creates powerful search algorithms.
@@ 2.2 Genetic Algorithms and Programs
Genetic algorithms (GA), a specific evolutionary algorithm, are often used for optimization tasks where the solution can be encoded as a sequence of values. GAs search the solution space efficiently and have been applied in domains such as reinforcement learning (whiteson_evolutionary_2006) and prompt optimization (secheresse_gaapo_2025). Genetic programming, which evolves tree-structured representations, has specifically been used for Automatic Heuristic Design to automatically generate and adapt decision-making rules tailored to specific tasks (burke_automatic_2007). These processes have advanced with the introduction of LLM-based evolutionary program search, where large language models guide the exploration, refinement, and assessment of complex programs in place of these less generalizable, rule-based frameworks (zhang_understanding_2024; shum_hybrid_2025). In this direction, LLMs have also been employed as mutation operators, leveraging their semantic understanding to generate more meaningful code modifications (lehman_evolution_2022). Evolutionary approaches have also been applied to neural architecture search (wistuba_survey_2019) and quality-diversity optimization through algorithms like MAP-Elites (mouret_illuminating_2015). More recently, evolutionary prompt engineering has been used as a method to optimize prompts for specific tasks, treating prompts as evolvable entities subject to selection and crossover (zhou_large_2023). Finally, Mind Evolution (lee_evolving_2025) has explored inference-time genetic algorithm approach for reasoning tasks. While the research reports strong performance with specific models (e.g., 94.1% accuracy using Gemini 1.5 Pro on a trip planning benchmark), it exhibits substantial variation under alternative model choices (e.g., 48.1% with GPT-4o-Mini), highlighting sensitivity to model selection. These observations motivate our focus on improving robustness across model families in evolutionary search algorithms.
@@ 3 Methodology
We introduce MultiGA (Multi-Source Genetic Algorithm), a framework that harnesses the complementary strengths of a diverse pool of LLMs. Rather than relying on a single model for both generation and evaluation, MultiGA initializes its population using outputs from multiple LLMs, each contributing distinct variations shaped by its training and architecture. These candidate solutions are then iteratively recombined and refined using an independent evaluator model drawn from a different model family. This evaluator provides nuanced feedback and guides the selection of higher-quality offspring.
@@ 3.3 Evaluation and Selection
In order to guide the search toward progressively more accurate solutions, we apply a two-phase fitness evaluation function to every candidate in the population at each iteration of the algorithm. The first phase of the evaluation framework compares predicted labels against the ground truth to assign perfect scores when applicable, while the second phase employs an LLM-as-a-judge framework.
At generation $t$, the population is represented as $P_{t}=\{x_{1},x_{2},\dots,x_{n}\}$. Each candidate $x_{i}$ is assigned a fitness score by $E$, defined as
@@
where $s_{i}=1$ corresponds to a perfect solution and $s_{i}=0$ indicates an invalid one. In the second phase, the evaluator $E$ is provided with all relevant task information (e.g., the query, context, or constraints) and is instructed to judge the overall correctness and quality of the candidate. This approach ensures that scoring is based on semantic adequacy rather than superficial string similarity.
@@ 3.4 Recombination
To generate new candidates, we employ an LLM-based crossover operator using the same independent LLM $E$. Given two parent solutions, $E$ is provided with all task information along with the parent solutions and is prompted to synthesize a child that integrates the strengths of both. Unlike traditional token-level crossover rules, this operator leverages the generative capacity of LLMs to perform semantic recombination. The resulting offspring are diverse yet coherent, inheriting useful attributes from each parent while overcoming the constraints of rule-based recombination strategies.  [^1] An interactive demo of the MultiGA framework is available at https://anonymous.4open.science/w/multiga-8771/
@@ 4 Experiments
To assess whether seeding genetic algorithms with multiple LLMs enables deeper and more optimal exploration, we evaluate our framework across four tasks. Our aim is not to design highly specialized or task-specific algorithms, but rather to provide a general framework in which models configurations can be evaluated fairly.
@@
In all experiments, we used a consistent configuration across tasks to ensure comparability. We selected the top-3 ($k=3$) parents for recombination, each run was capped at a maximum of $T=5$ generations, we used a fitness threshold of $\tau=0.2$ for pruning low-quality candidates, and our target fitness was $\phi=1.00$ as the stopping criterion. The genetic algorithm itself is implemented as a general class that can be configured for any task using prompts defined in a task-specific file.
4.1 Model Selection
We select either GPT-5 or Qwen3-Next-80B (abbreviated as Qwen3) to serve as the evaluator and recombination model $E$, responsible for scoring candidate solutions and synthesizing offspring across generations. These models were selected to contrast proprietary and open-source frontier LLMs in the evaluator role.
A key design choice is that the generating LLMs $\mathcal{G}=\{g_{1},\dots,g_{m}\}$ used in each experiment are drawn from model families distinct from the evaluator $E$ to avoid potential bias. These models were selected to span diversity in size, architecture, and training objectives.
To evaluate the models, we define two sets of generating LLMs based on the evaluator used:
Evaluator: Qwen3
The generator set $\mathcal{G}_{\text{A}}$ includes: GPT-5, GPT-4o-mini, Llama-4-Maverick-17B, Phi-4-Mini, and Gemma-2-27b.
Evaluator: GPT-5
The generator set $\mathcal{G}_{\text{B}}$ includes: Llama-4-Maverick-17B, Phi-4-Mini, Gemma-2-27b, and Qwen3-Next-80B.
4.2 Ablation Study
As a point of comparison, we evaluate MultiGA against baselines in which the genetic algorithm process is identical, but the initial population is seeded exclusively from a single model. For all such comparisons, we use Qwen3 as the evaluator E, given its strong empirical performance in our experiments and accessibility. This design isolates the effect of diverse initialization while holding all other components fixed (see Table 2 2).
@@ 4.2 Ablation Study
$\mathcal{D}_{\texttt{SQL}}$ (BIRD Mini-Dev) 200 500 $\infty$
@@
$\mathcal{D}_{\texttt{BBQ}}$ (BBQ Bias Data) 207 6879 $\infty$
@@
Table 1: Datasets used in our experiments, with test set size, total available items, and number of label choices provided to the generating LLMs.
4.3 Experimental Datasets
Text-to-SQL ($\mathcal{D}_{\texttt{SQL}}$)
Text-to-SQL is a critical task for enabling non-technical users to query databases and extract insights without writing code. In industry, developing cost-efficient and accurate text-to-SQL frameworks is essential for building robust agentic systems. For this experiment, we used the BIRD mini-dev dataset, which contains 500 realistic industry-level questions paired with gold-standard SQL queries (li_can_2023). We partitioned the data into training and test sets, using the training set to construct positive and negative examples for each test query. Specifically, we embedded the training data and performed cosine similarity search to retrieve relevant natural language–SQL pairs for the current question. All prompting logic and retrieval configurations were defined in a task-specific configuration file to ensure reproducibility (see Appendix B).
Meeting Planning ($\mathcal{D}_{\texttt{NATPLAN}}$)
We next applied MultiGA to structured reasoning through the meeting scheduling benchmark introduced by zheng_natural_2024. In this task, the LLM must generate a plan that maximizes the number of valid meetings during a hypothetical trip to San Francisco. Solving it requires navigating multiple logistical constraints, such as travel distance between meeting locations, meeting durations, and participant availability. This benchmark thus provides a natural setting to test whether iterative recombination can improve complex structured outputs (see Appendix C). We note that the evaluation metric assesses only whether the generated output matches the number of valid meeting plans, rather than the exact plan itself. [^2] Meeting Planning Evaluation: https://github.com/google-deepmind/natural-plan
GPQA Science Questions ($\mathcal{D}_{\texttt{GPQA}}$)
To broaden the evaluation of MultiGA, we tested the framework on the GPQA benchmark of graduate-level science multiple-choice questions (rein_gpqa_2023). This dataset covers advanced scientific domains and provides a challenging benchmark beyond code generation and optimization tasks. We used the Diamond subset, which contains the highest-quality questions in the benchmark (see Appendix D). We randomized the answer choices in the prompt to avoid any positional bias that may arise in the LLM generators (ye2024justiceprejudicequantifyingbiases).
BBQ Bias Evaluation ($\mathcal{D}_{\texttt{BBQ}}$)
Finally, we evaluated whether MultiGA mitigates social bias using the BBQ benchmark (parrish_bbq_2022). We focused on the race and ethnicity subset, which contains over 6,000 examples. Each example presents an ambiguous or unambiguous scenario involving racial stereotypes, followed by a question designed to reveal whether the model exhibits biased behavior. Rather than using the benchmark in its standard multiple-choice format, we adapted it to an open-ended setting and subsequently extracted answers for accuracy evaluation (see Appendix E).
5 Results and Discussion
$\mathcal{D}_{\texttt{SQL}}$ $\mathcal{D}_{\texttt{NATPLAN}}$ $\mathcal{D}_{\texttt{GPQA}}$ $\mathcal{D}_{\texttt{BBQ}}$
Generation Model(s) Eval Model Acc Tok Acc Tok Acc Tok Acc Tok
Baselines
GPT-5 (0-shot) - 0.290 5942 0.775 10997 0.732 4500 0.951 128
GPT-5 (N-shot) - 0.480 6674 0.806 9873 - - - -
Qwen3 (0-shot) - 0.480 4624 0.796 7877 0.287 1418 0.961 168
Qwen3 (N-shot) - 0.490 5585 0.622 6345 - - - -
Single Seeding Model (Uses MultiGA Framework)
GPT-5 Qwen3 0.585 133085 0.950 81770 0.914 28148 0.961 15355
Llama-4 Qwen3 0.630 117428 0.750 107670 0.833 19859 0.995 23596
Gemma-2 Qwen3 0.475 68651 0.320 52673 0.696 22935 0.961 9523
Phi-4-Mini Qwen3 0.165 232313 0.000 - 0.646 13963 0.923 33629
MultiGA (Ensemble Configurations)
$\mathcal{G}_{\text{A}}$ Qwen3 0.705 157430 0.880 71499 0.959 13349 0.966 36574
$\mathcal{G}_{\text{B}}$ GPT-5 0.685 134930 0.690 92784 0.919 8304 0.874 37723
Table 2: Accuracy and token usage per question across tasks, comparing baselines (0-shot, NN-shot), individual seeding models, and ensemble configurations $\mathcal{G}_{\text{A}}$ and $\mathcal{G}_{\text{B}}$. $\mathcal{D}_{\texttt{SQL}}$ and $\mathcal{D}_{\texttt{NATPLAN}}$ use few-shot only (3-shot and 5-shot respectively). GPT-4o-mini replaces GPT-5 as $E$ for $\mathcal{D}_{\texttt{BBQ}}$ due to moderation guardrails (see §5.4).
Figure 2: These are various examples of response quality evolution across generations for the $\mathcal{G}_{\text{A}}$ variant across each task. Each dot represents a data point in the benchmark, and the score is the fitness score assigned by the evaluator LLM (in this setting Qwen3).
We evaluate task accuracy on four benchmarks; Text-to-SQL ($\mathcal{D}_{\texttt{SQL}}$), GPQA Science Questions ($\mathcal{D}_{\texttt{GPQA}}$), Meeting Planning ($\mathcal{D}_{\texttt{NATPLAN}}$), and BBQ Bias ($\mathcal{D}_{\texttt{BBQ}}$). For each task, we compare single-model seeding with multi-source seeding using $\mathcal{G}_{\text{A}}$ or $\mathcal{G}_{\text{B}}$, depending on the evaluator. Results are shown in Table 2.
Across all four benchmarks, we observe that seeding with $\mathcal{G}_{\text{A}}$ (Llama-4-Maverick-17B, Phi-4-Mini, Gemma-2-27b, GPT-4o-mini, GPT-5) and Qwen3 as $E$ yields consistently high accuracy while requiring only two paid-model calls (used for seeding the population of initial solutions) over the entire optimization process for each task.